Find two consecutive odd integers in which the sum is equal to one or more than three times the smallest
Answer by drj(1380)

Find two consecutive odd integers in which the sum is equal to one or more than three times the smallest.

Step 1. Let n be one odd integer and let n+2 be the next consecutive integer.

Step 2. Let 3n be three times the smallest.

Step 3. Then n+n+2=1+3n or 2n+2=1+3n. Then n=1 and n+2=3.

Step 4. ANSWER: The numbers are 1 and 3.

I hope the above steps were helpful.

For FREE Step-By-Step videos in Introduction to Algebra, please visit
http://www.FreedomUniversity.TV/courses/IntroAlgebra and for Trigonometry visit
http://www.FreedomUniversity.TV/courses/Trigonometry.

Good luck in your studies!

Respectfully,
Dr J